# NocPlace: Nocturnal Visual Place Recognition Using Generative and Inherited Knowledge Transfer

### 27 Feb 2024 v1

@@ 6 Conclusion
In this paper, we introduce a novel training and testing strategy tailored for VPR, aimed at tackling the challenges of cross-domain recognition between day and night scenarios. With the proposed unpaired day-night location dataset, we can transfer the appearance knowledge of nighttime to existing large-scale, multi-view VPR datasets. By minimizing the classification loss of the target domain and the domain discrepancy between the source and target domains, we further enhance the performance of night-to-day VPR. Extensive experiments have been conducted on comprehensive nighttime VPR datasets to substantiate our contributions. NocPlace, with its straightforward and scalable approach, not only addresses the degradation observed in existing VPRs during night-to-day recognition but will also benefit future VPR systems.
Supplementary Material
@@ 7 Datasets
7.1 NightCities
We supplement the NightCities with daytime instances and category directories, as shown in Fig. 5 and Fig. 6. Over the next 1-3 years, we will continue to maintain and expand this dataset.
Figure 5: Daytime examples of the NightCities dataset. We supplement the daytime examples of the NightCities dataset here. From left to right, top to bottom, they represent Tokyo, Paris, Zurich, Hong Kong, Washington, San Francisco, Sydney, Munich, and Vancouver.
@@ 7.1 NightCities
7.2 SF-XL-NC
We trained NEG-CUT on the NightCities dataset and used it to process a subset of SF-XL, resulting in SF-XL-NC. As shown in Fig. 7, we present some examples demonstrating our success in coupling nighttime knowledge with other desired knowledge for VPR.
@@ 7.2 SF-XL-NC
8 Experiments
8.1 Comparing I2I translation methods
Method FID $\downarrow$ L2 $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ Avg Time $\downarrow$ CycleGAN 63.93 117.59 12.88 0.68 10 ms CUT 63.20 129.47 11.12 0.60 3.2 ms NEGCUT 63.54 121.62 11.75 0.60 3.2 ms ESGDE - - - - 62000 ms
Table 6: Comparsion of different unpaired I2I methods
We trained a range of state-of-the-art unpaired I2I generation methods to the NightStreets dataset, as detailed in Tab. 6. Interestingly, the disparities in outcomes among these methods are considerably less pronounced than those observed on other public datasets, presenting a deviation from some established research trends. However, one aspect is clear: the increasingly popular Diffusion Model, though capable of generating high-quality images, is notably time-intensive. For example, processing a training set of 100,000 VPR images with ESGDE [70] on a 3090 Ti GPU requires approximately 72 days. Given that our primary research focus remains on VPR, we did not extensively dwell on these issues and adopted NEG-CUT [57] for our study.
8.2 Qualitative results
Some qualitative results are shown in Fig 8 and Fig. 9. The figure allows to understand the strengths of NocPlace in a more visual and intuitive way.
Figure 9: Qualitative results on the Tokyo 24/7 dataset. Each column represents a query (first rows) and the first predicted result from the database. We can see that NocPlace handles extreme lighting variations better than previous methods.

### 21 Mar 2024 v2

@@ 6 Conclusion
In this paper, we introduce a novel training and testing strategy designed for VPR to tackle the challenges of cross-domain retrieval and localization between day and night scenarios. With the proposed unpaired day-night urban scene dataset, we transfer the appearance knowledge of nighttime to an existing large-scale daytime VPR dataset. By minimizing the classification loss in the target domain and the domain discrepancy between source and target domains, we further enhance the performance of night-to-day VPR. Considering practical use, a retrieval strategy based on partial divide-and-conquer was proposed for the first time. Extensive experiments conducted on comprehensive nighttime VPR test sets substantiate our contributions. As a straightforward and scalable approach, NocPlace addresses the degradation observed in existing VPRs during night-to-day recognition and holds promise for improving future VPRs and other downstream tasks in computer vision.
7 Overview
The table of contents of the supplementary content is as follows.
  • Discussions
    • The importance of Night-to-Day VPR
    • The novelty of generative knowledge transfer
    • The choice of backbones and methods
    • The choice of test sets
  • Experiments
    • Comparisons with SoTA Methods (2024.03)
    • Comparing I2I translation methods
    • Qualitative results
  • Datasets
    • The proposed datasets
    • The Datasets of VPR
8 Discussions
8.1 The importance of Night-to-Day VPR.
Night-to-day VPR, as a sub-problem of VPR, may make readers think it is unimportant. We provide the following insights:
  • Many studies [78, 37, 15, 4, 79] have shown that nighttime problems are equally important as daytime problems. For practical, we do not want intelligent robots and AR glasses to be available only during the day.
  • While LiDAR is a nighttime alternative to cameras, it costs much more than the latter, is much larger, and does not provide color information.
  • The nighttime problem differs other fundamental problems (e.g., visual scaling, occlusion) in that it can be predicted when it will occur and decomposed from the parent problem. However, researchers always try to solve it in independent systems. The divide-and-conquer strategy used here is novel.
8.2 The novelty of generative knowledge transfer.
The simplicity of generative knowledge transfer (GKT) may lead some readers to question its novelty. We provide the following explanations:
  • ToDay GAN [4] is proposed to generate daytime images for the retrieval phase of VPR, while NocPlace reverses the idea of ToDay GAN and generates night images for the training phase of VPR. If a study like NocPlace existed, it would inevitably cite To-Day GAN. While researching NocPlace, we searched Google Scholar many times and found that no one had implemented this simple and seemingly uninnovative idea.
  • DSA (ICCV’2023) [46] is a recent contemporaneous work with NocPlace, which also utilizes an image generation network to obtain night-style training images. Experiments about DSA are supplemented in Tab. 6.
  • The fact that some experienced researchers  [^3] https://github.com/AAnoosheh/ToDayGAN/issues/10 consider the idea of GKT to be infeasible suggests that it has not been successfully implemented or proven effective before.
  • Generative knowledge transfer is straightforward, but its implementation relies on high-quality images and advanced image generation methods. We solved the first problem and made many attempts at the latter.
8.3 The choice of backbones and methods.
Although the reasons for choosing backbones and methods have been stated in the paper, readers may still have concerns. We provide the following additional explanations:
  • The scaling law (OpenAI, 2020) [33] is a basic law of deep learning. As shown in previous experiments, improving the model architecture or increasing the descriptor dimension can correspondingly improve the results, demonstrating the scalability of NocPlace. Experienced researchers can foresee the impact of foundation models. Moreover, foundation models for VPR can be studied as an independent research problem (e.g., AnyLoc, CricaVPR, and SALAD).
  • We highly recognize the performance of AnyLoc, CricalVPR, and SALAD in public data sets and the inspiration they bring to the community. However, VPR is one of the technologies used in robots and AR glasses, which have stringent requirements for power, computing, and latency caused by limited resources (NeuroGPR [82], Science Robotics’2023).
  • NocPlace belongs to the one-stage VPR method, so there is no comparison experiment with the two-stage VPR method. In fact, NocPlace outperforms some two-stage state-of-the-art methods.
8.4 The choice of test sets.
Although the experiments in this article cover many test sets, two test sets are still missing. We provide the following explanations:
  • Gardens Point [24] (2014) . The official link is no longer available to obtain the Gardens Point dataset, and there is a community version  [^4] https://github.com/aghagol/loop-detection/issues/2. However, this dataset has become unchallenging due to its size (200 daytime and 200 nighttime images).
  • Alderley [45] (2012). The official link is no longer available to obtain the Alderley dataset  [^5] https://open.qcr.ai/dataset/alderley/, and we did not find any community versions.
9 Experiments
9.1 Comparing with SoTA Methods (2024)
We give the following observations:
  • On Tokyo 24/7 and MSLS Night, the Recall@1 of NocPlace is higher than the Recall@1 of CricaVPR (CVPR’2024) and AnyLoc (RAL’2024).
  • NocPlace is far better than DSA (ICCV’2023) and EigenPlaces (ICCV’2023) on all data sets.
  • SALAD (CVPR’2024) achieves the best results on the three data sets.
  • CricaVPR, SALAD, and AnyLoc all use ViT-based DINO-v2 [48] as feature extractors, and their feature dimensions are much higher than NocPlace.
  • The parameters shown in Tab. 6 are the parameters of the backbone. NocPlace is the model with the smallest parameters and the fastest feature extraction speed.
Table 6: Recall@1 / 5 / 10 on Night-to-Day VPR datasets.
Method Backbone
Feat.
Dim.
Params $\downarrow$
Speed $\uparrow$
(it/s)
SVOX
Night
Tokyo 24/7
MSLS
Night
SF-XL
Night
DSA [46]
DSA [46]
EigenPlaces [11]
NocPlace
VGG-16
Res-101
Res-50
Res-50
512
2048
2048
2048
138
45
24
24
58.86
63.30
198.90
198.90
8.3 / 19.0 / 23.1
10.4 / 20.0 / 25.3
58.8 / 76.9 / 82.6
75.6 / 87.5 / 91.4
61.3 / 74.9 / 80.6
69.5 / 83.5 / 87.3
93.0 / 96.2 / 97.5
95.2 / 96.8 / 98.1
3.6 / 3.6 / 7.3
3.6 / 5.5 / 9.1
3.6 / 10.9 / 10.9
69.1 / 76.4 / 80.0
23.6 / 30.9 / 34.5
28.3 / 35.8 / 39.7
AnyLoc [35]
SALAD [32]
CricaVPR [18]
DINOv2-g
DINOv2-B
DINOv2-B
49152
8448
4096
1100
86
86
3.24
30.40
56.24
77.5 / 90.6 / 94.5
97.0 / 99.0 / 99.1
85.1 / 95.0 / 96.7
91.7 / 97.5 / 98.7
97.5 / 98.4 / 98.4
93.0 / 97.5 / 98.1
3.6 / 5.5 / 16.4
90.9 / 100.0 / 100.0
50.9 / 78.2 / 81.8
SALAD-Noc
DINOv2-B
8448
86
30.40
96.4 / 99.0 / 99.1
97.5 / 98.4 / 98.7
94.9 / 98.2 / 100.0
9.2 Comparing I2I translation methods
Table 7: Comparsion of different unpaired I2I methods
Method FID $\downarrow$ L2 $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ Avg Time $\downarrow$
CycleGAN [89] 63.93 117.59 12.88 0.68 10 ms
CUT [50] 63.20 129.47 11.12 0.60 3.2 ms
NEGCUT [73] 63.54 121.62 11.75 0.60 3.2 ms
ESGDE [88] - - - - 62000 ms
We trained a range of state-of-the-art unpaired I2I translation methods on NightStreets, as detailed in Tab. 7. Interestingly, the disparities in outcomes among these methods are considerably less pronounced than those observed on other public datasets. However, one aspect is clear: the increasingly popular Diffusion Model, though capable of generating high-quality images, is notably time-intensive. For example, processing a training set of 100,000 VPR images with ESGDE [88] on a 3090 Ti GPU requires approximately 72 days. Given that our primary research focuses on VPR, we did not extensively dwell on these issues and adopted NEG-CUT [73] for our study.
9.3 Qualitative results
Some qualitative results are shown in Fig 5 and Fig. 6. The figure allows us to understand the strengths of NocPlace more visually and intuitively.
Figure 6: Qualitative results on the Tokyo 24/7 dataset. Each column represents a query (first row) and the first predicted result from the database. We can see that NocPlace handles extreme lighting variations better than previous methods.
@@ 10 Datasets
10.1 The proposed datasets
10.1.1 NightCities
We supplement the NightCities with daytime images and category directories, as shown in Fig. 7 and Fig. 8.
Figure 7: Daytime examples of the NightCities dataset. We supplement the daytime images of the NightCities dataset here. From left to right, top to bottom, they represent Tokyo, Paris, Zurich, Hong Kong, Washington, San Francisco, Sydney, Munich, and Vancouver.
@@ 10.1.1 NightCities
10.1.2 SF-XL-NC
@@
We trained NEG-CUT [73] on the NightCities dataset and used it to process a subset of SF-XL [8], resulting in SF-XL-NC. As shown in Fig. 9, we present some examples demonstrating our success in coupling nighttime knowledge with other desired knowledge for VPR.
10.2 Datasets of VPR
We test all models on a large number of night test sets, which helps to thoroughly understand each method’s strength and weaknesses. Below is a short description for each of the datasets.
10.2.1 SF-XL Night
[8] is a huge dataset covering the whole city of San Francisco with over 41M images. Its test set covers the same with a less dense set of 2.8M images. SF-XL night queries [6] are used which downloaded from Flickr for the area of San Francisco.
10.2.2 Tokyo 24/7
[5] is a challenging dataset from the center of Tokyo. The database is made from GoogleStreetView, whereas the queries are a collection of smartphone photos from 105 places, and each place is photographed during the day, at sunset and at night. This results in 315 queries, each to be geolocalized independently.
10.2.3 MSLS Night
[74] is the Mapillary Street-Level Sequences dataset, which has been created for image and sequence-based VPR. The dataset consists of more than 1M images from multiple cities, although only a small subset is used for evaluation.
10.2.4 SVOX Night
[12] is a cross-domain VPR dataset collected in a variety of weather and lighting conditions. It includes a large-scale database sourced from GoogleStreetView images spanning the city of Oxford. The night queries are extracted from the Oxford RobotCar dataset [43].
10.2.5 Aachen v1.1
[58] depicts the old inner city of Aachen, Germany. The database images used to build the reference scene representations were all taken during daytime with hand-held cameras over a period of about two years. The datasets offers query images taken at daytime and at nighttime. All query images were taken with mobile phone cameras, i.e., the Aachen Day-Night dataset considers the scenario of localization using mobile devices, e.g., for Augmented or Mixed Reality. The nighttime query images were taken using software HDR to create (relatively) high-quality images that are well-illuminated.
10.2.6 RobotCar Seasons v2
[58] depicts the city of Oxford, UK. The reference and query images were captured by three synchronized cameras mounted on a car, pointing to the rear-left, rear, and rear-right, respectively. The images were recorded by driving the same route over a period of 12 months. One traversal is used to define a reference condition and the reference scene representation. Other traversals, covering different seasonal and illumination conditions, are used for query. All images were recorded in sequences. The RobotCar Seasons dataset represents an autonomous driving scenario, where it is necessary to localize images taken under varying seasonal conditions against a (possibly outdated) reference scene representation. Compared to the Aachen Day-Night dataset, the nighttime images of the RobotCar Seasons dataset exhibit significantly more motion blur and are of lower image quality.